Write a quadratic equation having the given numbers as solutions. -4 and -2
Type in standard form.

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

Write a quadratic equation having the given numbers as solutions. -4 and -2
Type in standard form.
---------
(x+4)*(x+2) = 0

Answer by jim_thompson5910(35256)

x = -4 or x = -2

x+4=0 or x+2=0

(x+4)(x+2)=0

x^2+2x+4x+8=0

x^2+6x+8=0